OpenStudy (anonymous):

i am needing help 3/10 and 6/7 fractions with common denominators then the sum

OpenStudy (akshay_budhkar):

ok the basic concept in such questions is that u take the LCM of the two denominators

OpenStudy (akshay_budhkar):

so in this case we take the LCM of 7 and 10 , which is 70, so we multiply the numerator and denominator with the numbers so the we get equivalent fractions with denominators 70

OpenStudy (akshay_budhkar):

so for the first fraction we get\[{3 \over 10} = {3 \over 10} \times {7\over 7} = {21 \over 70}\]

OpenStudy (akshay_budhkar):

for the next fraction we have\[{6 \over 7} = {6\over7} \times { 10 \over 10} = { 60 \over 70}\]

OpenStudy (akshay_budhkar):

so now you can add them together\[{21 \over 70 } + {60 \over 70 } = { { 21+60} \over 70 } = {81 \over 70 }\]

OpenStudy (akshay_budhkar):

is it clear?

OpenStudy (anonymous):

yes ma'am it is my daughter didnt believe me so i asked for help lol

OpenStudy (akshay_budhkar):

* sir :D hehe no problem it is always better to confirm :D

OpenStudy (anonymous):

sorry

OpenStudy (akshay_budhkar):

no problem, ppl generally think my name is that of a girl lol :D And welcome to openstudy :D

OpenStudy (anonymous):

thanks i will be using this more often with her